faith1001234:

The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?

RinaRinaBoBina:

I will help you on the first part. the first equation is y = -4x + 4 y = 4x + c 2= 4(-8) + c 2 = -32 + c 34 = c y = 4x + 34